Willie can do the job in 6 days. Myra can do the same job in 4.5 days. If they work together, how long will it take to complete the job?
Answer by nerdybill(7384)

Willie can do the job in 6 days. Myra can do the same job in 4.5 days. If they work together, how long will it take to complete the job?
.
Let x = time (days) for both
then
x(1/6 + 1/4.5) = 1
multiply both sides by 6(4.5):
x(4.5 + 6) = 6(4.5)
x(10.5) = 27
x = 2.57 days
or
x = 2 days 13 hours and 43 minutes